# D-69120 Heidelberg, Germany

### v1

@@ 7 Discussion and Caveats
We also assume perfect alignment of the grain rotation axis with the direction of the radiation i.e. $\theta=0$. This is based on the fact, that grains preferentially align with the direction of incoming radiation in the absence of magnetic fields (see LH07). Though in general magnetic fields might be present and the grains may not reach suprathermal angular velocities. For a more detailed picture, one would need DDSCAT simulations sampling the whole range of $\theta$-angles. These can then be used to calculate phase space trajectories of grains in $\theta$ and $\omega$, possibly including magnetic fields. Such an approach may allow to verify if the assumption of fixed orientation was justified, and if not, to gauge how large the deviation is. However, this is beyond the scope of our current study.
@@
We have shown, that aggregated grains have a different functional form of $Q_{\Gamma}(\lambda)$ than compact grains and therefore do not reach as high angular velocities as compact grains would in the same circumstances, with differences of up to an order of magnitude for the smallest grains. However we see that our results and the canonical parametrization agree for the least dense and largest grains (BA, $250\,\mathrm{nm}$). If the trend of increasing $c$ for larger grain size and lower monomer-density (see Fig. 12) holds true, we would expect $\Gamma_{\mathrm{RAT}}$ to further increase for less dense and larger grains. That lower monomer-density leads to higher torques makes intuitive sense considering it increases the total surface area of the grain and therefore the cross section for interaction with radiation just as increasing the grain size does. While the torque increases, $\omega_{\mathrm{RAT}}$ however does not increase further for larger grains and in some cases even decreases again for the largest grains. Aggregate grains larger than the ones simulated here may potentially not reach much higher angular velocities, as the drag also increases with size.
The dust grains were assumed to have no relative velocity to the gas, but it is possible, that grains close to strong radiation fields experiencing linear acceleration by the radiation pressure will reach high drift velocities relative to the surrounding gas. The rotation of the grain would then also be influenced by mechanical torques (Reissl et al. 2022) working in a similar way to RATs, with gas collisions instead of photon collisions.
@@ 8 Summary and Outlook
In this paper, we present a numerical study of the impact of grain shape on the radiate torque (RAT) alignment efficiency. This particular alignment mechanism predicts a sensitivity of the grain alignment efficiency with respect to the magnitude of the radiation field as well as the materials and shape of the dust grains. As a source of radiation we consider O-type stars, an active galactic nuclei (AGN) and the interstellar radiation field (ISRF) typical for the Milky Way. The dust grains are modelled by means of ballistic aggregation and migration (BAM) of monomers of various sizes. The aggregation process is biased to control the size and number of connections of the grain aggregates. We calculate the optical dust grain properties in the discrete dipole approximation using the DDCAT code considering the refractive indices of pure carbonaceous and silicate grains, respectively. Given the optical properties of the dust aggregates and the radiation energy density we calculate the maximal angular velocity $\omega_{\mathrm{RAT}}$ according to RAT alignment theory assuming stable alignment. Our major findings are as follows:
  • We report that the ISRF produced average angular velocities $\omega_{\mathrm{RAT}}=10^{6}\,\mathrm{s}^{-1}$ for both carbon and silicate grains where as the O-type stars and AGN sources produced much larger $\omega_{\mathrm{RAT}}$ in a range of $10^{9}\,\mathrm{s}^{-1}$ to $10^{11}\,\mathrm{s}^{-1}$ for carbon grains and $10^{8}\,\mathrm{s}^{-1}$ to $10^{10}\,\mathrm{s}^{-1}$ for silicate grains.
@@
  • Considering an entire ensemble of BAM grain aggregates, we propose a new parametrization for the RAT efficiency that follows a broken power law with a constant part for shorter wavelengths followed by two distinct power laws to cover longer wavelengths. The proposed parametrization describes the average angular velocities $\omega_{\mathrm{RAT}}$ reached by the ensembles of grains more precisely, than the canonical parametrization derived for compact grains.
  • According to our DDSCAT calculations individual grain aggregates may have sign changes in the RAT efficiency over the entire wavelength regime reducing the net torque integrated over wavelength. However, the appearance of such sign changes does impact the average grain spin-up behaviour of the entire ensemble only marginally.
Our results for a parametrization with two power-laws for grains composed of monomers, potentially improved by expanding to a parametrization for the distribution of $\omega_{\mathrm{RAT}}$ values, can make detailed studies on grain rotation including grain size, porosity, and material easier by replacing computationally expensive DDSCAT simulations of individual grains. The assumption of unobscured source radiation may be dropped by including effects of the surrounding gas and dust around the grain to better gauge the effect of RATs further away from the radiation source or in dense environments. Furthermore, the results of this work may be further constrained by including a general rotation of the grains, possibly including magnetic fields, where individual grains are followed along their pathways in the rotation phase-space.
@@ Appendix A Two power-law parametrization fit parameters
For completeness we provide the full set of fitted parameters of the two-power-law RAT parametrization in table 4 for CA grains and in table 5 for SC grains.
monomer-density $q_{1}$ $q_{2}$ $b$ $a_{\mathrm{eff}}$ [nm] $c$ $a$
BA $-6.690\pm 0.007$ $-1.0032\pm 0.0003$ $14\pm 5$ 50 $0.071\pm 0.004$ $3.202\pm 0.026$
100 $0.091\pm 0.004$ $3.083\pm 0.022$
150 $0.154\pm 0.007$ $2.853\pm 0.021$
200 $0.265\pm 0.012$ $2.629\pm 0.018$
250 $0.295\pm 0.013$ $2.588\pm 0.018$
BAM1 $-6.659\pm 0.011$ $-1.02855\pm 0.00022$ $12\pm 4$ 50 $0.0315\pm 0.0027$ $3.21\pm 0.04$
100 $0.064\pm 0.003$ $2.891\pm 0.025$
150 $0.112\pm 0.005$ $2.655\pm 0.018$
200 $0.159\pm 0.007$ $2.521\pm 0.018$
250 $0.200\pm 0.009$ $2.434\pm 0.017$
BAM2 $-6.073\pm 0.007$ $-1.01848\pm 0.00018$ $13\pm 5$ 50 $0.0232\pm 0.0014$ $3.13\pm 0.03$
100 $0.0358\pm 0.0016$ $2.916\pm 0.023$
150 $0.070\pm 0.003$ $2.614\pm 0.022$
200 $0.096\pm 0.004$ $2.481\pm 0.019$
250 $0.099\pm 0.005$ $2.467\pm 0.019$
Table 4: Parameters for CA
monomer-density $q_{1}$ $q_{2}$ $b$ $a_{\mathrm{eff}}$ [nm] $c$ $a$
BA $-6.168\pm 0.008$ $-0.7434\pm 0.0004$ $13\pm 5$ 50 $0.062\pm 0.004$ $3.11\pm 0.03$
100 $0.123\pm 0.007$ $2.785\pm 0.026$
150 $0.227\pm 0.011$ $2.523\pm 0.020$
200 $0.299\pm 0.015$ $2.412\pm 0.020$
250 $0.389\pm 0.017$ $2.311\pm 0.017$
BAM1 $-6.393\pm 0.007$ $-0.7033\pm 0.0003$ $12\pm 4$ 50 $0.0360\pm 0.0022$ $3.18\pm 0.03$
100 $0.075\pm 0.005$ $2.832\pm 0.029$
150 $0.169\pm 0.009$ $2.494\pm 0.021$
200 $0.201\pm 0.010$ $2.428\pm 0.020$
250 $0.224\pm 0.011$ $2.387\pm 0.020$
BAM2 $-5.386\pm 0.008$ $-0.8477\pm 0.0003$ $11\pm 5$ 50 $0.0266\pm 0.0014$ $2.779\pm 0.027$
100 $0.065\pm 0.004$ $2.36\pm 0.03$
150 $0.082\pm 0.005$ $2.256\pm 0.028$
200 $0.112\pm 0.007$ $2.127\pm 0.023$
250 $0.128\pm 0.006$ $2.078\pm 0.020$
Table 5: Parameters for SC
Acknowledgements.
Special thanks go to Thiem Hoang for numerous enlightening discussion. S.R., J.A.J., and R.S.K. acknowledge financial support from the Heidelberg cluster of excellence (EXC 2181 - 390900948) “STRUCTURES: A unifying approach to emergent phenomena in the physical world, mathematics, and complex data”, specifically via the exploratory project EP 4.4. S.R. and R.S.K. also thank for support from Deutsche Forschungsgemeinschaft (DFG) via the Collaborative Research Center (SFB 881, Project-ID 138713538) ’The Milky Way System’ (subprojects A01, A06, B01, B02, and B08). And we thank for funding form the European Research Council in the ERC synergy grant “ECOGAL – Understanding our Galactic ecosystem: From the disk of the Milky Way to the formation sites of stars and planets” (project ID 855130). The project made use of computing resources provided by The Länd through bwHPC and by DFG through grant INST 35/1134-1 FUGG. Data are in part stored at SDS@hd supported by the Ministry of Science, Research and the Arts and by DFG through grant INST 35/1314-1 FUGG.

### v2

@@ 7 Discussion and Caveats
We also assume perfect alignment of the grain rotation axis with the direction of the radiation i.e. $\Theta=0$. This is based on the fact, that grains preferentially align with the direction of incoming radiation in the absence of magnetic fields (see LH07). Though magnetic fields are present in the typical environments probed here and the grains may not reach suprathermal angular velocities. For a more detailed picture, one would need DDSCAT simulations sampling the whole range of $\Theta$-angles and, subsequently, be used to calculate phase space trajectories of grains in $\Theta$ and $\omega$, as well as including typical magnetic fields. Such an approach may allow to verify if the assumption of fixed orientation was justified, and if not, to gauge how large the deviation is. However, this is beyond the scope of our current study.
@@
We have shown, that aggregated grains have a different functional form of $Q_{\Gamma}(\lambda)$ than compact grains and therefore do not reach as high angular velocities as compact grains would in the same circumstances, with differences of up to an order of magnitude for the smallest grains. However we see that our results and the canonical parametrization agree for the least dense and largest grains (BA, $250\,\mathrm{nm}$). If the trend of increasing $c$ for larger grain size and lower monomer-density (see Fig. 12) holds true, we would expect $\Gamma_{\mathrm{RAT}}$ to further increase for less dense and larger grains. The relation of lower monomer-density leading to higher torques makes intuitive sense considering it increases the total surface area of the grain and therefore the cross section for interaction with radiation just as increasing the grain size does. While the torque increases, $\omega_{\mathrm{RAT}}$ however does not increase further for larger grains and in some cases even decreases again for the largest grains. Aggregate grains larger than the ones simulated here may potentially not reach much higher angular velocities, as the drag also increases with size.
In environments with highly ionized gas and charged grains, such as an AGN, plasma drag in addition to gas drag and photon drag might additionally play a role for grain dynamics (Draine & Lazarian 1998). However, the charging of BA dust adds an additional level of complexity where individual monomers may carry different charges and even undergo some stochastic fluctuation over time (see e.g. Matthews et al. 2013; Ma et al. 2013; Matthews et al. 2018). Including the plasma drag in the context of BA goes beyond the scope of this paper. However, the difference of the total grain charge of a BA compared to an equivalent sphere should be less than one order of magnitude as shown in Ma et al. (2013) and Matthews et al. (2018). Hence, we estimate that the canonical parametrization of RAT would still overestimate the resulting angular velocity $\omega_{\mathrm{RAT}}$ for ballistic aggregates in proximity of an AGN. The dust grains were assumed to have no relative velocity to the gas, but it is possible, that grains close to strong radiation fields experiencing linear acceleration by the radiation pressure will reach high drift velocities relative to the surrounding gas. We emphasize that the resulting rotation of the BA grain would then also be influenced by mechanical torques (METs, see e.g. Lazarian & Hoang 2007b; Das & Weingartner 2016; Hoang et al. 2018; Reissl et al. 2023) working in a similar way to RATs, with gas collisions instead of photon collisions. However, the question of how METs may enhance the grain rotation in the presence of RATs is still a matter of debate (see also Lazarian & Hoang 2021).
We emphasize once again that our modeling of RATs does not include the dynamical processes associated with the presence of an external magnetic field. As noted in LH07 for an analytical RAT model, the appearance of attractors in the $\Theta-\omega$ phase space is tightly connected to the ratio ${q=\max(Q_{\mathrm{e1}})/\max(Q_{\mathrm{e2}})}$ when an external field is present. Later, this finding was confirmed by Herranen et al. (2019) and Herranen (2020) based on grain shapes modeled as Gaussian ellipsoids. We note, however, that a similar parameter $q$ for pure mechanical alignment was not found (Das & Weingartner 2016; Reissl et al. 2023). Hence, it remains to be seen if the ratio $q$ is a reliable proxy to predict the long-term stable alignment with the direction of the magnetic field, considering RATs as a spin-up mechanism of ballistic aggregates. Moreover, in the presence of a magnetic field, this is not only a matter of the efficiency of the RATs. Due to internal dissipation processes in paramagnetic and ferromagnetic materials, a fraction of the angular velocity $\omega$ is transferred into internal heat (see e.g. Davis & Greenstein 1951; Dolginov & Mytrophanov 1976; Purcell 1979). Consequently, the $\omega$ reported in our study may be lower, depending on the magnetic field strength and grain orientation. A similar effect of a reduction in $\omega$ for METs is shown by Reissl et al. (2023). However, to what degree a magnetic field would reduce the maximal possible angular velocity $\omega_{\mathrm{RAT}}$ of BAM aggregates remains to be addressed in forthcoming publications.
@@ 8 Summary and Outlook
In this paper, we present a numerical study of the impact of grain shape on the radiative torque (RAT) efficiency and its dependence on the magnitude of the radiation field as well as the material and porosity of the dust grains. As a source of radiation we consider O-type stars, an active galactic nuclei (AGN) and the interstellar radiation field (ISRF) typical for the Milky Way. The dust grains are modeled by means of ballistic aggregation and migration (BAM) of monomers of various sizes. The aggregation process is biased to control the size and number of connections of the grain aggregates. We calculate the optical dust grain properties in the discrete dipole approximation using the DDSCAT code considering the refractive indices of pure carbonaceous and silicate grains, respectively. Given the optical properties of the dust aggregates and the radiation energy density we calculate the maximal angular velocity $\omega_{\mathrm{RAT}}$ according to RAT alignment theory. We note again that the conclusions presented here are still preliminary in nature, as stable alignment of grain rotation axis and direction of incoming radiation was assumed and no magnetic fields were included. Our major findings are as follows:
  • We report that the ISRF produced average angular velocities $\omega_{\mathrm{RAT}}=10^{6}\,\mathrm{s}^{-1}$ for both carbon and silicate grains where as the O-type stars and AGN sources produced much larger $\omega_{\mathrm{RAT}}$ in a range of $10^{8}\,\mathrm{s}^{-1}$ to $10^{10}\,\mathrm{s}^{-1}$ for carbon grains and $10^{7}\,\mathrm{s}^{-1}$ to $10^{9}\,\mathrm{s}^{-1}$ for silicate grains.
@@
  • Considering an entire ensemble of BAM grain aggregates, we propose a new parametrization for the RAT efficiency that follows a broken power-law with a constant part for shorter wavelengths followed by two distinct power-laws to cover longer wavelengths. The proposed parametrization describes the average angular velocities $\omega_{\mathrm{RAT}}$ reached by the ensembles of grains more precisely, than the canonical parametrization derived for compact grains.
  • According to our DDSCAT calculations individual grain aggregates may have sign changes in the RAT efficiency over the entire wavelength regime reducing the net torque integrated over wavelength. However, the appearance of such sign changes does impact the average grain spin-up behavior of the entire ensemble only marginally.
  • The stronger stellar source and the AGN produce angular velocities potentially high enough for rotational disruption of the grains. The results for the angular velocities obtained here could be used to study the rotational disruption of ballistic aggregate grains on the level of individual grains in the future.
Our results for a parametrization with two power-laws for grains composed of monomers, potentially improved by expanding to a parametrization for the distribution of $\omega_{\mathrm{RAT}}$ values, can make detailed studies on grain rotation including grain size, porosity, and material easier by replacing computationally expensive DDSCAT simulations of individual grains. Further improvements could be made by dropping the assumption of unobscured source radiation, including isotropic radiation, surface effects, $\mathrm{H}_{2}$-formation and other additional drag effects. Especially in order to apply the results to grain alignment a general orientation of the grains as well as magnetic fields should be added to complement the analysis.
Acknowledgements.
We thank the anonymous referee for a constructive report. Special thanks go to Thiem Hoang for numerous enlightening discussions. R.S.K. also thanks the Harvard-Smithsonian Center for Astrophysics and the Radcliffe Institute for Advanced Studies for their hospitality during his sabbatical, and the 2024/25 Class of Radcliffe Fellows for highly interesting and stimulating discussions. S.R., J.A.J., and R.S.K. acknowledge financial support from the Heidelberg cluster of excellence (EXC 2181 - 390900948) “STRUCTURES: A unifying approach to emergent phenomena in the physical world, mathematics, and complex data”, specifically via the exploratory project EP 4.4. S.R. and R.S.K. also thank for support from Deutsche Forschungsgemeinschaft (DFG) via the Collaborative Research Center (SFB 881, Project-ID 138713538) ’The Milky Way System’ (subprojects A01, A06, B01, B02, and B08). And we thank for funding form the European Research Council in the ERC synergy grant “ECOGAL – Understanding our Galactic ecosystem: From the disk of the Milky Way to the formation sites of stars and planets” (project ID 855130). The project made use of computing resources provided by The Länd through bwHPC and by DFG through grant INST 35/1134-1 FUGG. Data are in part stored at SDS@hd supported by the Ministry of Science, Research and the Arts and by DFG through grant INST 35/1314-1 FUGG.
@@ Appendix A Two power-law parametrization fit parameters
For completeness we provide the full set of fitted parameters of the two-power-law RAT parametrization in table A for CA grains and in table A for SC grains.
[…]
tableFit parameters for CA monomer-density $q_{1}$ $q_{2}$ $b$ $a_{\mathrm{eff}}$ [nm] $c$ $a$ BA $-6.690\pm 0.007$ $-1.0032\pm 0.0003$ $14\pm 5$ 50 $0.071\pm 0.004$ $3.202\pm 0.026$ 100 $0.091\pm 0.004$ $3.083\pm 0.022$ 150 $0.154\pm 0.007$ $2.853\pm 0.021$ 200 $0.265\pm 0.012$ $2.629\pm 0.018$ 250 $0.295\pm 0.013$ $2.588\pm 0.018$ BAM1 $-6.659\pm 0.011$ $-1.02855\pm 0.00022$ $12\pm 4$ 50 $0.0315\pm 0.0027$ $3.21\pm 0.04$ 100 $0.064\pm 0.003$ $2.891\pm 0.025$ 150 $0.112\pm 0.005$ $2.655\pm 0.018$ 200 $0.159\pm 0.007$ $2.521\pm 0.018$ 250 $0.200\pm 0.009$ $2.434\pm 0.017$ BAM2 $-6.073\pm 0.007$ $-1.01848\pm 0.00018$ $13\pm 5$ 50 $0.0232\pm 0.0014$ $3.13\pm 0.03$ 100 $0.0358\pm 0.0016$ $2.916\pm 0.023$ 150 $0.070\pm 0.003$ $2.614\pm 0.022$ 200 $0.096\pm 0.004$ $2.481\pm 0.019$ 250 $0.099\pm 0.005$ $2.467\pm 0.019$
[…]
tableFit parameters for SC monomer-density $q_{1}$ $q_{2}$ $b$ $a_{\mathrm{eff}}$ [nm] $c$ $a$ BA $-6.168\pm 0.008$ $-0.7434\pm 0.0004$ $13\pm 5$ 50 $0.062\pm 0.004$ $3.11\pm 0.03$ 100 $0.123\pm 0.007$ $2.785\pm 0.026$ 150 $0.227\pm 0.011$ $2.523\pm 0.020$ 200 $0.299\pm 0.015$ $2.412\pm 0.020$ 250 $0.389\pm 0.017$ $2.311\pm 0.017$ BAM1 $-6.393\pm 0.007$ $-0.7033\pm 0.0003$ $12\pm 4$ 50 $0.0360\pm 0.0022$ $3.18\pm 0.03$ 100 $0.075\pm 0.005$ $2.832\pm 0.029$ 150 $0.169\pm 0.009$ $2.494\pm 0.021$ 200 $0.201\pm 0.010$ $2.428\pm 0.020$ 250 $0.224\pm 0.011$ $2.387\pm 0.020$ BAM2 $-5.386\pm 0.008$ $-0.8477\pm 0.0003$ $11\pm 5$ 50 $0.0266\pm 0.0014$ $2.779\pm 0.027$ 100 $0.065\pm 0.004$ $2.36\pm 0.03$ 150 $0.082\pm 0.005$ $2.256\pm 0.028$ 200 $0.112\pm 0.007$ $2.127\pm 0.023$ 250 $0.128\pm 0.006$ $2.078\pm 0.020$